# Hybrid-Integrated InGaAs/InP SPAD Arrays for Quantum Communications

### v1

@@ 1 Introduction
Here, we report the development of linear arrays of InGaAs/InP SPADs and their application in a hybrid-integrated QKD receiver. Figure 1 gives an overview of the work. We use a quasi-planar coupling (QPC) technique (Figure 1.a) that allows us to efficiently couple the arrays to silica waveguide chips. We demonstrate the first GHz-gating of an InGaAs/InP SPAD array and utilise an electronic circuit that enables efficient collective driving as well as individual pixel bias optimisation to overcome non-uniform gain (Figure 1.b). Characterisation of our SPAD devices yields competitive performance: on all pixels a four device array exhibits 15% detection efficiency, dark count rates per pixel below 8 kHz, and afterpulsing below 4% at a 100 ns deadtime. For the two devices used in the hybrid QKD receiver we achieve a mean pixel dark count rate of 1.96 kHz, or $1.93\times 10^{-6}$ Hz/gate. Further, we find that GHz-gating significantly suppresses the inter-pixel crosstalk. Finally, to demonstrate applicability to quantum communications, we present results from a time-bin encoded BB84 QKD experiment (Figure 1.c). The setup consists of a chip-based transmitter and our hybrid integrated receiver, which combines a low-loss passive silica decoding circuit with an InGaAs/InP SPAD array. The QKD system achieves a positive secure key rate over 100 km of fibre spool and, using room-temperature detectors, we exceed 2 Mbps secure key rate at 0 dB channel attenuation. These results demonstrate a viable path toward compact and practical QKD receivers suitable for metropolitan and access quantum networks.
@@ 2.1 GHz-Gated InGaAs/InP SPAD Arrays
Our SPADs were operated in gated-mode at 1 GHz with ’self-differencing circuits’ used to filter the output signals [30], an approach that has been widely applied to QKD [31, 32]. However, the above approach has traditionally been applied to single detectors, and so had to be adapted for array operation. Figure 1.b shows a conceptual diagram of our supporting circuit; the devices are driven collectively through a shared cathode by the same 1 GHz gate signal, yet retain individual readout channels via separate anodes. Further, the DC bias of each pixel can be adjusted individually. This architecture has three advantages. First, the expensive driving circuit systems (signal generator, amplifier, high voltage DC source) are shared between the detectors, making scaling to larger channel numbers economical. Second, the timing synchronisation of gates between the detectors is implicit in the shared RF drive signal. Third, the adjustability of pixel biases is useful for both individual device optimisation and for QKD security due to the existence of the detector-efficiency mismatch attack [33, 34].
@@
Figure 2: Four-device SPAD array performance during 1 GHz gated-mode operation. a Dark count rate (DCR) and afterpulsing rate (APR) against single photon detection efficiency (SPDE) as the universal array bias is varied. Inset: SPDE against universal array bias across the most relevant voltage range. b Synchronous crosstalk probability across the array for each of the three victim pixels (at successively increasing distance) as the universal array bias was increased. c Array specificity: Detected count rates for the array as each device is illuminated, with dark count rates subtracted. All four devices have their DC bias set to achieve 15% SPDE. We see specificity of at least three orders of magnitude for all pixel combinations.
@@
The performance of a four-device linear SPAD array was characterised under 1550 nm laser illumination. The driving signal consisted of a 1 GHz square wave, with a 400 ps duty cycle of amplitude 12 Vpp, and a DC bias set close to the breakdown voltage (57 V). A 100 ns deadtime was applied to control the afterpulsing rate. Illumination was with a short ($\sim$ 30 ps) laser pulse with a repetition rate of 15.625 MHz, such that one in every 64 gates were illuminated. The laser was projected onto the SPAD apertures using a 50x microscope lens with a numerical aperture of 0.42. Each laser pulse was calibrated to contain a mean flux of 0.2 photons, as measured in the plane of the SPAD aperture. Elevated counts in non-illuminated gates were used to calculate afterpulsing probabilities, whilst dark count rates were measured under no illumination. The results of the characterisation are shown in Figure 2a, demonstrating simultaneous operation across a four-device array with detection efficiencies above 15%, afterpulsing below 4%, and dark count rates per device below 8 kHz. The results represent the first reported GHz-gating of an InGaAs/InP SPAD array.
@@ 2.3 Quasi-Planar Coupling
To achieve optical coupling between our silica chips and linear SPAD arrays we used a QPC technique, illustrated in Figures 1.a and 1.b. At the output side of the chip an angled cut produces an angled facet. Light propagating horizontally within the waveguides undergoes total internal reflection and is projected downwards. The waveguides run a few microns below the surface of the chip, meaning that the light leaves the chip within a few microns of the tip. The QPC approach was chosen for several advantages. First, it is an established low-loss technique for coupling between silica fibre arrays and PICs. Second, the material silica has low propagation loss which is desirable for a QKD receiver. Third, the coupling orientation allows for the silica chip and SPAD array to be mounted in parallel planes, reducing the height of the final assembly.
With a straight-waveguide chip the QPC approach allowed us to conveniently couple the detectors to individual fibre inputs, whilst for the QKD receiver we could directly functionalise the silica PIC as part of the receiver optical circuit. The mode-field diameter of the waveguides in the silica chip was $\sim$10 µm, whilst the aperture diameter of our SPADs was 20 µm, leading to reasonable alignment tolerance in both horizontal and vertical axes. Alignment to arrays was achieved by simultaneously maximising the count rate of the devices at the extreme ends.
In Table 1 we provide data estimating the coupling loss of the QPC technique. The four-device SPAD array is operated with a balanced SPDE through a four-channel QPC chip, with known channel losses, and then compared to equivalent data with microscope illumination. For the four channels we find estimates of QPC-to-SPAD coupling losses ranging from 0.22 dB to 0.68 dB.
@@
System SPDE (QPC) / % 10.25 10.36 10.27 10.42
Waveguide Channel Loss / dB 1.97 0.72 0.89 1.15
@@
QPC-to-SPAD Coupling Loss / dB 0.22 0.68 0.40 0.22
Table 1: Estimation of optical coupling loss across a four-device SPAD array using quasi-planar coupling (QPC). ’System SPDE’ is the efficiency as measured from the input fibre connector of the waveguide chip. The individual pixel biases have been adjusted to achieve an approximately uniform system SPDE of 10% despite variations in waveguide channel loss. By comparing the system SPDE with the known SPAD SPDE, and after compensating for the waveguide chip channel loss, we can estimate the optical coupling loss between the waveguide chip and the SPAD using the QPC technique.
@@ 2.4 QKD Receiver Assembly
Our QKD receiver waveguide chip contained an asymmetric Mach Zehnder interferometer (AMZI) structure with a 500 ps arm length difference and the two outputs of the AMZI proceeding to a QPC angled facet. An optical insertion loss of 1.8 dB was measured at 1550 nm for the waveguide chip. For the QKD receiver we used two detectors from a three-device array (characterisation in Supplemental Document S2). The three-device array was used due to lower dark counts than the four-device array. The waveguide chip was mechanically held against the SPAD array, rather than permanently attached with glue, to allow us to use both samples for future experiments. The entire assembly was mounted on a cold plate set to -30°C and with the SPAD array further temperature stabilised to within $\pm$0.01°C by a dynamically controlled thermo-electric cooler. The assembly was held within a nitrogen-purged atmosphere to prevent condensation during cooling.
@@ 3 BB84 QKD Experiment
To reinforce the suitability of the SPAD array for quantum communications we performed a chip-based BB84 QKD experiment (Figure 1.c ). In the transmitter we utilise a previously published edge-coupled hybrid PIC encoder [18]. Within Alice, a 1 GHz train of gain-switched laser pulses was filtered with passband width 0.1 nm and centered at 1552.8 nm. Polarisers and polarisation controllers were used to control the propagation of polarisation throughout the system. A VOA at the output of Alice was used to bring the photon flux down to the single photon regime. The communication channel either consisted of a VOA to emulate fibre loss or a real fibre spool. In Bob, an external phase modulator was used to perform active measurement basis switching. The phase modulator added 2.4 dB to the receiver circuit, bringing the total Bob loss to 4.2 dB. The active components of the system were controlled using field programmable gate arrays (FPGAs) adapted from a previously published standalone chip-based QKD system[7].
@@
Figure 3: Plot of raw bit rate, QBER and secure bit rate against channel attenuation with detectors at -30°C. The loss of a VOA in the quantum channel was gradually increased to emulate increasing fibre channel distance. Equivalent fibre channel distance is calculated at 0.18 dB/km. Data was also taken over a real 100 km fibre spool (stars), with a loss of 19.2 dB (elevated due to fibre connector losses). The dotted lines show the results of simulations based on measured system parameters.
@@ 4 Discussion
We have demonstrated a QKD receiver with hybrid-integrated non-cryogenic detectors. The results highlight a realistic pathway toward the manufacture of compact QKD receivers with non-cryogenic cooling requirements. In classical fibre-optic communication, small-form-factor pluggable transceivers, typically only a few centimetres in size, integrate the full photonic stack of lasers, modulators, and detectors. In contrast, quantum networks have relied on external detector systems, precluding similar miniaturisation. Our work suggests a viable route toward achieving this goal.
@@
Acknowledgment We thank Robert Woodward for supplying BB84 finite key analysis code and guidance. We thank M. Sanzaro for productive conversations and ideas. J.A.D. thanks R.V. Penty for his guidance and supervision. J.A.D. acknowledges funding from the UK’s Engineering and Physical Sciences Research Council under the Industrial Cooperative Awards in Science & Technology (CASE) programme.

### v2

@@ 1 Introduction
Here, we report the development of linear arrays of InGaAs/InP SPADs and their application in a hybrid-integrated QKD receiver. Figure 1 gives an overview of the work. We use a quasi-planar coupling technique (Figure 1.a) that allows us to efficiently couple the arrays to silica waveguide chips[26]. We demonstrate the first GHz-gating of an InGaAs/InP SPAD array and utilise an electronic circuit that enables efficient collective driving as well as individual pixel bias optimisation to overcome non-uniform SPDE (Figure 1.b). Characterisation of our SPAD devices yields competitive performance: on all pixels a four device array exhibits 15% detection efficiency, dark count rates per pixel below 8 kHz, and afterpulsing below 4% at a 100 ns deadtime. For the two devices used in the hybrid QKD receiver we achieve a mean pixel dark count rate of 1.96 kHz, or $1.93\times 10^{-6}$ Hz/gate. Further, we find that GHz-gating significantly suppresses the inter-pixel crosstalk. Finally, to demonstrate applicability to quantum communications, we present results from a time-bin encoded BB84 QKD experiment (Figure 1.c). The setup consists of a chip-based transmitter and our hybrid integrated receiver, which combines a low-loss passive silica decoding circuit with an InGaAs/InP SPAD array. The QKD system achieves a positive secure key rate over 100 km of fibre spool and, using room-temperature detectors, we exceed 2 Mbps secure key rate at 0 dB channel attenuation. These results demonstrate a viable path toward compact and practical QKD receivers suitable for metropolitan and access quantum networks.
@@ 2.1 GHz-Gated InGaAs/InP SPAD Arrays
Our SPADs were operated in gated-mode at 1 GHz with ’self-differencing circuits’ used to filter the output signals [31], an approach that has been widely applied to QKD [32, 33]. However, the above approach has traditionally been applied to single detectors. Figure 1.b shows a conceptual diagram of our supporting circuit; the devices are driven collectively through a shared cathode by the same 1 GHz gate signal, yet retain individual readout channels via separate anodes. Each readout channel used its own self-differencing filter and the DC bias of each pixel could be adjusted individually through a small bias on the anode. Otherwise, no modifications were necessary compared to a conventional self-differencing SPAD circuit. This architecture has three advantages. First, the expensive driving circuit systems (signal generator, amplifier, high voltage DC source) are shared between the detectors, making scaling to larger channel numbers economical. Second, the timing synchronisation of gates between the detectors is implicit in the shared RF drive signal. Third, the adjustability of pixel biases is useful for both individual device optimisation and for QKD security due to the existence of the detector-efficiency mismatch attack [34, 35].
@@
Figure 2: Four-device SPAD array performance during 1 GHz gated-mode operation. a Dark count rate and afterpulsing probability (after 100 ns deadtime) against single photon detection efficiency (SPDE) as the universal array bias is varied. Inset: SPDE against universal array bias across the most relevant voltage range. b Synchronous crosstalk probability across the array for each of the three victim pixels (at successively increasing distance) as the universal array bias was increased. c Array specificity: Detected count rates for the array as each device is illuminated, with dark count rates subtracted. All four devices have their DC bias set to achieve 15% SPDE. We see specificity of at least three orders of magnitude for all pixel combinations.
@@
The performance of a four-device linear SPAD array was characterised under 1550 nm laser illumination. The driving signal consisted of a 1 GHz square wave, with a 400 ps duty cycle of amplitude 12 Vpp, and a DC bias set close to the breakdown voltage ($\sim$57 V). A 100 ns deadtime was applied to control the afterpulsing rate. Illumination was with a short ($\sim$30 ps) laser pulse with a repetition rate of 15.625 MHz, such that one in every 64 gates were illuminated. The laser was projected onto the SPAD apertures using a 50x microscope lens with a numerical aperture of 0.42. Each laser pulse was calibrated to contain a mean flux of 0.2 photons, as measured in the plane of the SPAD aperture. Elevated counts in non-illuminated gates were used to calculate afterpulsing probabilities, whilst dark count rates were measured under no illumination. The results of the characterisation are shown in Figure 2a, demonstrating simultaneous operation across a four-device array with detection efficiencies above 15%, afterpulsing below 4%, and dark count rates per device below 8 kHz. The results represent the first reported GHz-gating of an InGaAs/InP SPAD array.
@@ 2.3 Quasi-Planar Coupling
Our silica chips were manufactured using an ion-exchange method[46, 47], capable of achieving a propagation loss of < 0.2 dB/cm. A mode-field diameter of 10 µm was used with a minimum bend radius of 1 mm. To achieve optical coupling between our silica chips and linear SPAD arrays we used a quasi-planar coupling technique, illustrated in Figures 1.a and 1.b. At the output side of the chip an angled cut produces an angled facet. Light propagating horizontally within the waveguides undergoes total internal reflection and is projected downwards. The waveguides run a few microns below the surface of the chip, meaning that the light leaves the chip within a few microns of the tip. The quasi-planar coupling approach was chosen for several advantages. First, it is an established low-loss technique for coupling between silica fibre arrays and PICs. Second, the material silica has low propagation loss which is desirable for a QKD receiver. Third, the coupling orientation allows for the silica chip and SPAD array to be mounted in parallel planes, reducing the height of the final assembly.
With a straight-waveguide chip the quasi-planar coupling approach allowed us to conveniently couple the detectors to individual fibre inputs, whilst for the QKD receiver we could directly functionalise the silica PIC as part of the receiver optical circuit. As the aperture diameter of our SPADs was 20 µm, whilst the mode-field diameter was 10 µm, this led to reasonable alignment tolerance in both horizontal and vertical axes. Alignment to arrays was achieved by simultaneously maximising the count rate of the devices at the extreme ends.
In Table 1 we provide data estimating the coupling loss of the quasi-planar coupling technique. The four-device SPAD array is operated with a balanced SPDE through a four-channel quasi-planar coupling chip, with known channel losses, and then compared to equivalent data with microscope illumination. For the four channels we find estimates of silica-to-SPAD coupling losses ranging from 0.22 dB to 0.68 dB.
@@
System SPDE (Silica) / % 10.25 10.36 10.27 10.42
Silica Chip Channel Loss / dB 1.97 0.72 0.89 1.15
@@
Silica-to-SPAD Coupling Loss / dB 0.22 0.68 0.40 0.22
Table 1: Estimation of optical coupling loss across a four-device SPAD array using quasi-planar coupling. ’System SPDE’ is the efficiency as measured from the input fibre connectors of the silica chip. The individual pixel biases have been adjusted to achieve an approximately uniform system SPDE of 10%, despite variations in silica chip channel loss. By comparing the system SPDE with the known SPAD SPDE, and after compensating for the silica chip channel loss, we can estimate the optical coupling loss between the silica chip and the SPAD using the quasi-planar coupling technique.
@@ 2.4 QKD Receiver Assembly
Our QKD receiver waveguide chip contained an asymmetric Mach Zehnder interferometer structure with a 500 ps arm length difference and the two outputs of the interferometer proceeding to a quasi-planar coupling angled facet. An optical insertion loss of 1.8 dB was measured at 1550 nm for the waveguide chip. For the QKD receiver we used two detectors from a three-device array (characterisation in Supplemental Document S2). The three-device array was used due to lower dark counts than the four-device array. The waveguide chip was mechanically held against the SPAD array, rather than permanently attached with glue, to allow us to use both samples for future experiments. The entire assembly was mounted on a cold plate set to -30°C and with the SPAD array further temperature stabilised to within $\pm$0.01°C by a dynamically controlled thermo-electric cooler. The assembly was held within a nitrogen-purged atmosphere to prevent condensation during cooling.
@@ 3 BB84 QKD Experiment
To reinforce the suitability of the SPAD array for quantum communications we performed a chip-based BB84 QKD experiment (Figure 1.c ). In the transmitter we utilise a previously published edge-coupled hybrid PIC encoder [18]. Within Alice, a 1 GHz train of gain-switched laser pulses was filtered with passband width 0.1 nm and centered at 1552.8 nm. Polarisers and polarisation controllers were used to control the propagation of polarisation throughout the system. A variable optical attenuator at the output of Alice was used to bring the photon flux down to the single photon regime. The communication channel either consisted of a variable optical attenuator to emulate fibre loss or a real fibre spool. In Bob, an external phase modulator was used to perform active measurement basis switching. The phase modulator added 2.4 dB to the receiver circuit, bringing the total Bob loss to 4.2 dB. The active components of the system were controlled using field programmable gate arrays (FPGAs) adapted from a previously published standalone chip-based QKD system[7].
@@
Figure 3: Plot of raw bit rate, QBER and secure bit rate against channel attenuation with detectors at -30°C. The loss of a variable optical attenuator in the quantum channel was gradually increased to emulate increasing fibre channel distance. Equivalent fibre channel distance is calculated at 0.18 dB/km. Data was also taken over a real 100 km fibre spool (stars), with a loss of 19.2 dB (elevated due to fibre connector losses). The dotted lines show the results of simulations based on measured system parameters.
@@ 4 Discussion
We have demonstrated a QKD receiver with hybrid-integrated non-cryogenic detectors. The results highlight a realistic pathway toward the manufacture of compact QKD receivers with non-cryogenic cooling requirements. It remains the case that the highest performance detectors for QKD are superconducting nanowire single photon detectors (SNSPDs), which provide superior detection efficiency (which increases secure bit rate) and lower dark count rate (which increases range). However, the significant size, weight, power and cost of SNSPDs systems seems sure to limit them to the most cost insensitive of applications. The integrated SPAD-array-based devices demonstrated within this work are best-suited to fulfilling the large-scale deployment of QKD in metropolitan and access quantum networks, where the high-volume production capability and cost efficiency are prioritised over secure bit rate and range. In classical fibre-optic communication, small-form-factor pluggable transceivers, typically only a few centimetres in size, integrate the full photonic stack of lasers, modulators, and detectors. In contrast, quantum networks have relied on external detector systems, precluding similar miniaturisation. Our work suggests a viable route toward achieving this goal.
@@
Acknowledgment We thank Robert Woodward for supplying BB84 finite key analysis code and guidance. We thank M. Sanzaro for productive conversations and ideas. J.A.D. thanks R.V. Penty for his guidance and supervision. J.A.D. acknowledges PhD funding from the UK’s Engineering and Physical Sciences Research Council under the Industrial Cooperative Awards in Science & Technology (CASE) programme.